4x-4/3y=-2
-16+16/3y=8
the given system of linear equations has one solution, no solution, or an infinite number of solutions

Answer by lwsshak3(11628)

4x-4/3y=-2
-16+16/3y=8
the given system of linear equations has one solution, no solution, or an infinite number of solutions
***
..
4x-4y/3=-2
-16+16y/3=8
..
4y/3=4x+2
16y/3=24
..
4y=12x+6
16y=72
..
16y=48x+24
16y=72
subtract
48x-72+24=0
48x=48
x=1
y=72/16=9/2
point of intersection: (1, 9/2)
given system of linear equations has one solution(ans.)